OpenStudy (anonymous):

Help! Explain as well! Identify the correct slope and y intercept of the equation 2x + 4y = 12. Answer slope = one half; y intercept at (0, 4) slope = one half; y intercept at (0, 3) slope = -negative one half; y intercept at (0, 3) slope = 2; y intercept at (0, 12)

OpenStudy (anonymous):

The first thing to do is to make the relation explicitly as function of y: \[2x+4y=12 \rightarrow 4y = 12-2x\] Dividing by 4 gives you \[y=3-1/2x\] So the slope is 1/2, and the y intercept can be found by solving y = 0

OpenStudy (anonymous):

Sorry, by solving for x = 0!

OpenStudy (anonymous):

Which gives y = 3 =)

OpenStudy (anonymous):

So, to summarize: 2x+4y=12 has the slope -1/2 and y intercept (0.3)

OpenStudy (anonymous):

Hmm.... I still don't understand...

hero (hero):

y = mx + b m = slope b = y -intercept If you write the given equation in that form, you will be able to find the slope and y-intercept